OpenStudy (anonymous):

MEDAL WILL BE GIVIN!!! What are the equivilant fraction and decimal for 33 1/3%? A. 1/33 and 0.3 B. 1/33 and 3 C. 3 1/3 and 0.33 D 1/3 and 0.33

rebeccaxhawaii (rebeccaxhawaii):

divide 1 by 3 so that will be your answer

OpenStudy (anonymous):

. . . 3

rebeccaxhawaii (rebeccaxhawaii):

so its D

OpenStudy (anonymous):

can you help me with more?

rebeccaxhawaii (rebeccaxhawaii):

sure

OpenStudy (anonymous):

ok thx, What are the equivalent fraction and decimal for 20% A. 1/20 and 0.2 B. 1/5 and 0.2 C. 1/50 and 2.2 D. 1/5 and 2

rebeccaxhawaii (rebeccaxhawaii):

well divide 1 by what do you get

OpenStudy (anonymous):

so its C?

rebeccaxhawaii (rebeccaxhawaii):

no not quite

OpenStudy (anonymous):

. . . but what do I divide then?

OpenStudy (anonymous):

its D?

rebeccaxhawaii (rebeccaxhawaii):

so divide the first number by the second and if you get the decimal and if it equals the decimal on the side thats your answer

OpenStudy (anonymous):

im really confused, wouldn't that make 0??

rebeccaxhawaii (rebeccaxhawaii):

no it will be a decimal

rebeccaxhawaii (rebeccaxhawaii):

it will be a decimal less than 1

OpenStudy (anonymous):

0.2?

OpenStudy (anonymous):

so its A

OpenStudy (anonymous):

wright??

rebeccaxhawaii (rebeccaxhawaii):

oh wait sorry

rebeccaxhawaii (rebeccaxhawaii):

i read the problem wrong

OpenStudy (anonymous):

. . .

OpenStudy (anonymous):

so can you help me still?

rebeccaxhawaii (rebeccaxhawaii):

it would be b bc 20 times 5 is 100 %

OpenStudy (anonymous):

k, so does that mean num 1 is wrong?

rebeccaxhawaii (rebeccaxhawaii):

yes

OpenStudy (anonymous):

no cuz its a fraction

OpenStudy (anonymous):

wait what? whats wrong with 1?

OpenStudy (anonymous):

y isnt anyone else saying anything?

OpenStudy (pawanyadav):

B is right one

OpenStudy (anonymous):

oh, for 1 or 2?

OpenStudy (pawanyadav):

2

OpenStudy (anonymous):

k, @rebeccaxhawaii

rebeccaxhawaii (rebeccaxhawaii):

what

OpenStudy (pawanyadav):

Because 20%=20/100=1/5 And 1/5=0.2

OpenStudy (anonymous):

another one?? Estimate 78% of 20. A. 14 B. 16 C. 19 D. 2 i think its C but not sure

rebeccaxhawaii (rebeccaxhawaii):

no its not c

OpenStudy (pawanyadav):

78×20/100=78/5~16 So B

rebeccaxhawaii (rebeccaxhawaii):

divide 20 by 4 so its B

rebeccaxhawaii (rebeccaxhawaii):

its would be 15 if it was 75% so 78% would be 16

OpenStudy (anonymous):

oh ok thx again

rebeccaxhawaii (rebeccaxhawaii):

yup

OpenStudy (anonymous):

i have to go i fanned u rebecca its ok if we message later?

rebeccaxhawaii (rebeccaxhawaii):

yes of course

OpenStudy (anonymous):

kk i will fan u too @Pawanyadav

OpenStudy (anonymous):

bye

OpenStudy (pawanyadav):

Bye